40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sun Sep 29 15:58:19 2019, gautam, Update, LSC, POX locking attempt POXlockAttempt.png
    Reply  Mon Sep 30 14:16:43 2019, gautam, Update, LSC, POX PD checkout - solved 
Message ID: 14915     Entry time: Mon Sep 30 14:16:43 2019     In reply to: 14910
Author: gautam 
Type: Update 
Category: LSC 
Subject: POX PD checkout - solved 

I confirmed that there is light incident on the POX photodiode. So the problem must lie downstream in the demod / whitening / AA electronics. With the PRM aligned (i.e. PRFPMI config with all DoFs uncontrolled), I could see the flashing beam on an IR card. I could also see the spikes in DC power incident on the photodiode using the "DC Monitor" port on the photodiode head and an oscilloscope.

Update 245 pm: I confirmed that I could see a 11 MHz sine wave by connecting the POX11 RFPD output cable at the 1Y2 end to an oscilloscope. The amplitude of this signal was also changing, corresponding to the cavity fringing in and out of resonance. I couldn't, however, see any signal on the RFPDmon port, or the I/Q demodulated output ports. So as of now, the culprit seems to be something on the Demod board. Further investigations underway...

Update 315pm: I did the following checks:

  1. Checked the LO signal level into a 50ohm input scope - it was ~720 mVpp, which was compatible with the LO level into the POY Demod board, so the LO signal level couldn't be to blame.
  2. Connected an RF funcgen to the PD input of the demod board. Drove it at 11.066210 MHz, 50 mVpp, and saw a signal 400 cts-pp in the CDS system - so the demod + digitizaiton electronics also seemed fine.
  3. #2, coupled with the fact I could see no signal at the RF-mon port of the demod board (even though there was a signal visible at the cable coming to 1Y2) suggested that the cable routing the POX11 PD output from the Heliax-breakout in 1Y2 to the demod board was busted - indeed this was the case!
  4. Koji replaced the cable without changing its length, and now the XARM locks readily 👏 . I ran ASS and got TRX ~ 0.95. See Attachment #1
Quote:

Look for the POX beam with an IR viewer.

ELOG V3.1.3-